Minutes of the 30th SOHO SWT Meeting

Max-Planck-Institut für Aeronomie, Lindau, Germany 23 March 2000

2

Contents

1 Welcoming Address 5 2 Action Items 5 3 Agree Agenda and Action Revision 5 4 Announcements 5 5 Spacecraft status 6 6 Ground System Status 6 7 Instrument status 6

7.1 GOLF (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.2 VIRGO (C. Frölich) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.3 MDI (R. Bush) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.4 SUMER . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.5 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.6 EIT (J.-P. Delaboudiniere) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 7.7 UVCS (J. Kohl) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 7.8 LASCO (R. Schwenn) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 7.9 SWAN (J.L. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 7.10 CELIAS (P. Bochsler) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 7.11 COSTEP (R. Mueller-Mellin) . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 7.12 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

8 Status & Demonstration of SOHO Archive 7 9 MEDOC Campaigns 8 10 Future PR and Outreach Activities 8 11 JOPs and Collaborations 8 12 Strategic Planning 8 13 Workshops and Meetings 9 14 AOB 9 Annex 1: Agenda 11 Annex 2: List of Participants 13 Annex 3: Spacecraft Status 15 Annex 4: SOHO Ground System Status 29 Annex 5-1: GOLF Status 35 Annex 5-2: VIRGO Status 39 Annex 5-3: MDI Status 41 Annex 5-4: SUMER Status & Future plans 51 Annex 5-5: SUMER Report on Azimuth Pointing Mechanism 57 Annex 5-6: CDS Status 67 Annex 5-7: EIT Status 71 Annex 5-8: UVCS Status 73 Annex 5-9: LASCO Status 81 Annex 5-10: SWAN Status 83 Annex 5-11: CELIAS Status 87 Annex 5-12: COSTEP Status 91 Annex 5-13: ERNE Status 99 Annex 6: Data Issues and Status of the new SOHO Archive 103 Annex 7: SOHO Archive Development Status at MEDOC 113 Annex 8: MEDOC 2000 Campaigns 115 Annex 9: Planned and Ongoing PR - Outreach 125 Annex 10: Joint Observing Programs & Collaborations 137 Annex 11: SOHO JOP Proposal Guidelines 141 Annex 12: Strategic Planning 145 Annex 13: SOHO Meetings & Events 153

1 Welcoming Address

B. Fleck welcomed the attendees. He noted the high number of participants, and the fact that representatives for all instruments were present.

K. Wilhelm welcomed the attendees to Max-Planck-Institut für Aeronomie on behalf of Director S. Solanki.

2 Action Items

Action 30-1: On PIs: Provide L. Sánchez Duarte^{1} with default data policy for archival data, if any restrictions are desired.

Action 30-2: On PIs: Provide updated information for the SOHO Bibliography & Publications database^{2}, with particular emphasis on Ph.D. Theses.

Action 30-3: On PIs: Request team members to send paper copies of papers in journals (other than ApJ and A&A) to the SOHO PS office^{3}.

3 Agree Agenda and Action Revision

No changes to the agenda (see annex 1). List of attendees - annex 2.

Actions Revision

Action 29-1: Closed On PIs: Appoint contact persons for planning of joint SOHO-HESSI observations.

Action 29-2: Ongoing/Closed On SOCs: Collect and coordinate plans for joint SOHO-HESSI observations.

Action 28-1: De facto Closed On European PIs: All the instrument teams are required to check if their flight software and ground software is Y2K compatible.

4 Announcements

J. Bruner will step down as SOHO Programme Manager. NASA has issued a vacancy notice for her replacement.

There will be an ϵ Flight Recertification Review at Goddard Space Flight Center on 13 April 2000.

A NASA Senior Review of SOHO will be held next spring.

1lsanchez@esa.nascom.nasa.gov

2Submit information to soho@esa.nascom.nasa.gov

3SOHO PS Office, NASA GSFC, Mail Code 682.3, SOHO EAF, Bldg 26 Room G-1, Greenbelt, MD 20771, USA

5 Spacecraft status

B. Fleck presented the spacecraft status (annex 3) on behalf of H. Schweitzer and M. Chaloupy, including the main events since SWT-29, ACU software patches, spacecraft subsystem status, and statistics on spacecraft anomalies.

6 Ground System Status

B. Fleck presented the ground system status (annex 4) on behalf of the FOT Manager E. Nace. The FOT Technical Manager is H. Benefield. The Observatory Engineers are B. Sapper, K. Hibbard and N. Piston.

He thanked the PIs on behalf of the FOT for their continued support. K. Wilhelm returned the thanks on behalf of the PIs.

7 Instrument status

The PIs presented the status of their instruments (see annex 5).

7.1 GOLF (A. Gabriel)

XXX (see annex 5-1)

7.2 VIRGO (C. Frölich)

(see annex 5-2)

7.3 MDI (R. Bush)

(see annex 5-3)

7.4 SUMER

K. Wilhelm presented the status of SUMER’s calibration, publications, detectors, and future plans (annex 5-4). U. Schühle presented a report on the SUMER azimuth pointing mechanism (annex 5-5).

7.5 CDS (R. Harrison)

(see annex 5-6)

7.6 EIT (J.-P. Delaboudiniere)

(see annex 5-7)

7.7 UVCS (J. Kohl)

(see annex 5-8)

7.8 LASCO (R. Schwenn)

R. Schwenn presented LASCO’s status on behalf of R. Howard (see annex 5-9)

7.9 SWAN (J.L. Bertaux)

(see annex 5-10)

7.10 CELIAS (P. Bochsler)

(see annex 5-11)

7.11 COSTEP (R. Mueller-Mellin)

(see annex 5-12)

7.12 ERNE (J. Torsti)

(see annex 5-13)

8 Status & Demonstration of SOHO Archive

On behalf of L. Sanchez Duarte, B. Fleck presented some data issues and the status of the new SOHO archive (see annex 6).

The problem of wrong roll attitude values (telemetry only – not the physical roll attitude) reported in LZ CD-ROMS will most likely be solved without involving the CDHF, meaning that LZ CD-ROMS will not be changed. Instead, the on-line SOHO archive will contain the corrected data, including a time series indicating which spacecraft mode is in effect. This option was approved unanimously by the SWT.

The SOHO archive is nearing completion. Database definition and creation, catalogue extraction, and database population has been solved.

I. Scholl demonstrated the GUI for PIs to modify data access rights to individual users for individual data sets (see annex 7). Since most instruments now have an open data policy, this will be the default unless PIs provide alternate policies for their instruments to L. Sanchez Duarte (action 30-1).

Everyone is encouraged to try out the GUI for accessing the database^{4} and to provide feedback to the developers.

4http://www.medoc-ias.u-psud.fr/archive/

9 MEDOC Campaigns

J.-C. Vial presented results of MEDOC Campaigns 3 & 4, and plans for Campaigns 5 & 6 (see annex 8).

The proposed dates for the 6th campaign were, however, in conflict with SUMER plans for coordinated observations with TRACE. Possibilities for MEDOC campaigns in July and/or November/December should be investigated by J.-C. Vial.

10 Future PR and Outreach Activities

A summary of PR and outreach activities were presented by P. Brekke (see annex 9).

11 JOPs and Collaborations

S.V.H. Haugan presented the current status of ongoing work (see annex 10) with SOHO-HESSI collaborations, and a draft of the “JOP Cookbook” (annex 11). It was agreed that in exceptional circumstances, a Deputy JOP Leader may be present at the EOF for initial runs of JOPs. It was stressed that this must be a person with intimate knowledge of the observing programs and scientific objectives of the JOP, not just a “local contact” as required for subsequent runs. The use of a Deputy JOP Leader is to be decided on a case-by-case basis.

It was also pointed out that JOP Leaders should provide information on the publication record of earlier runs when asking for repeat runs of JOPs.

12 Strategic Planning

B. Fleck presented some viewpoints on strategic planning (see annex 12).

In a discussion on operational aspects, J. Gurman pointed out that most instrument teams are now operating on a slimmer budget than during the initial mission phase. This contributes to an efficient, but perhaps less “scientifically stimulating” atmosphere at monthly, weekly and daily planning meetings.

R. Harrison asked whether other instruments had strong objections against some changes to CDS planning that would involve more “pre-planned” periods, to ease the workload of CDS planners. No objections were raised, as long as flexibility to schedule JOPs would be maintained.

Web server statistics demonstrate that the real-time images, screensaver and Hot Shots have given a tremendous increase in traffic during the last months. Contributions from all instruments are strongly encouraged. Keep in mind, however, that some journals have strict rules with regard to results which has been publicised prior to submission.

Among other outreach activities, J. Kohl and R. Harrison mentioned more local projects. Funding and support for such activities is, however, difficult to obtain.

The PIs are requested to provide updated publication lists, and to request team members to send in copies of papers in journals other than ApJ and A&A (actions 30-2 and 30-3). The number of Ph.D. theses in the SOHO bibliography was spuriously low – the PIs should make sure that updated information is submitted.

Team members should also be reminded about the use of the standard acknowledgement phrase, “SOHO is a project of international cooperation between ESA and NASA”.

Attrition of people to other projects was raised as one concern, though this may be construed as a positive aspect when seen from the Solar community at large.

13 Workshops and Meetings

A list of known meetings can be found in annex 13. See the “Meetings” page^{5} on the SOHO Web server for updates.

Note the SOHO/ACE 2001 Symposium, “Solar and Galactic Composition”, 6-9 March 2000, in Bern (see annex 17 of SWT-29 meeting minutes for more information).

14 AOB

SWT-31 will be held in conjunction with the IAU meeting in Manchester, on Wednesday, 9 August 2000.

R. Bush invited to SWT-32 at Stanford, in conjunction with the AGU Fall Meeting, 15-19 December 2000.

5http://sohowww.nascom.nasa.gov/meetings/

10

Annex 1: Agenda

11

Annex 2: List of Participants

13

Attendees SOHO SWT-30

Name  Institute/Experiment  Address
Bertaux, J.L.  Service d’Aeronomie/SWAN  BP 3, Verrieres-Le-Buisson, F-91371 (bertaux@aerov.jussieu.fr)
Bochsler, P.  Univ. of Bern/CELIAS  Sidlerstr. 5, CH-3012 Bern (peter.bochsler@soho.unibe.ch)
Brekke, P.  ESA/SSD  NASA/GSFC, Code 682.3 (pbrekke@esa.nascom.nasa.gov)
Bush, R.  Stanford University/SOI-MDI  HEPL Annex B212, Stanford, CA 94305-4085 (rbush@solar)
Curdt, W.  MPAE  (curdt@linmpi.mpg.de)
Delaboudiniere, J.P. IAS/EIT  Univ. Paris XI, F-91405 Orsay Cedex (boudine@ias.fr)
Fleck, B.  ESA/SSD  NASA/GSFC, Code 682.3 (bfleck@esa.nascom.nasa.gov)
Fröhlich, C.  PMOD-WRC/VIRGO  CH-7260 Davos-Dorf
Gabriel, A.  IAS/GOLF  Bat 121, Univ. Paris XI, F-91405 Orsay Cedex (gabriel@ias.fr)
Germerott, D.  MPAE  (germerott@linmpi.mpg.de)
Gurman, J.  NASA/GSFC/EIT  (gurman@gsfc.nasa.gov)
Harrison, R.A.  RAL/CDS  RAL, Didcot, Oxon, OX11 OQX, UK

(r.harrison@rl.ac.uk)

Haugan, S.V.H.  ESA/SSD  NASA/GSFC Code 682.3 (shaugan@esa.nascom.nasa.gov)
Kjeldseth-Moe, O.  Univ. Oslo/CDS  Inst. Theor. Astrophys., Oslo, Norway (olavm@astro.uio.no)
Kohl, J.  SAO/UVCS  SAO, 60 Garden ST, Cambridge, MA 02138, USA

(jkohl@cfa.harvard.edu)

Landi, E.  MPAE/SUMER  (landi@linmpi.mpg.de)
Mueller-Mellin, R.  Univ. of Kiel/CEPAC  Inst. für Exp. u. Angewandte Physik, Univ. Kiel,

Leibnizstr. 19, D-24118 Kiel (mueller-mellin@kernphysik.uni-kiel.de)

Quemerais, E.  Service d’Aeronomie/SWAN  BP 3, Verrieres-Le-Buisson, F-91371 (quemerai@aerov.jussieu.fr)
Scholl, I.  IAS  IAS Bât 121, Univ. Paris XI, F 91405 ORSAY

(scholl@medoc-ias.u-psud.fr)

Schuehle, U.  MPAE/SUMER  MPAE Lindau (schuehle@linmpi.mpg.de)
Schwenn, R.  MPAE  D-37191 Katlenburg-Lindau

(schwenn@linmpi.mpg.de)

Solanki, S.  MPAE  D-37191 Katlenburg-Lindau
Stenborg, G.  MPAE  D-37191 Katlenburg-Lindau (stenborg@linmpi.mpg.de)
Torsti, J.  University of Turku/ERNE  Space Res. Lab., SF-20520 Turku (jarmo.torsti@utu.fi)
Vial, J.-C.  IAS/SUMER  vial@ias.u-psud.fr
Wenzel, K.P.  ESA/SCI-SO  ESTEC (kwenzel@estec.esa.nl)
Wilhelm, K.  MPAE/SUMER  MPAe Lindau (wilhelm@linmpi.mpg.de)

Annex 3: Spacecraft Status

15

Annex 4: SOHO Ground System Status

29

Annex 5-1: GOLF Status

35

Annex 5-2: VIRGO Status

39

Annex 5-3: MDI Status

41

Annex 5-4: SUMER Status & Future plans

51

Annex 5-5: SUMER Report on Azimuth Pointing Mechanism

57

Annex 5-6: CDS Status

67

Annex 5-7: EIT Status

71

Annex 5-8: UVCS Status

73

Annex 5-9: LASCO Status

81

Annex 5-10: SWAN Status

83 86

Annex 5-11: CELIAS Status

87 90

Annex 5-12: COSTEP Status

91

Annex 5-13: ERNE Status

99 102

Annex 6: Data Issues and Status of the new SOHO Archive

103

Annex 7: SOHO Archive Development Status at MEDOC

113

Annex 8: MEDOC 2000 Campaigns

115 124

Annex 9: Planned and Ongoing PR - Outreach

125

Annex 10: Joint Observing Programs & Collaborations

137 140

Annex 11: SOHO JOP Proposal Guidelines

141

Annex 12: Strategic Planning

145

Annex 13: SOHO Meetings & Events

153